A 200 seat theater charges $3.00 for adults and $1.50 for children. If all seats were filled and the total ticket income was $510 how many adults and how many childen were in the audience?
Answer by ewatrrr(10682)


Hi
full 200 seat theater
Let x and (200-x)represent the number of adult and student tickets respectively
Question states***
3x + 1.50(200-x) = $510
Solving for x
1.50x = 510 - 1.5*200
x = 210/1.5
x = 140, number of adult tickets. 60 the number of student tickets.
CHECKING our Answer***
$3*140 + $1.50*60 = $420 + $90 = $510